OpenStudy (anonymous):

Find cos if sin=-5/13

OpenStudy (cwrw238):



OpenStudy (anonymous):

so if cos is c=a/h, then . . . cos=-12/13?

OpenStudy (cwrw238):

yes

OpenStudy (anonymous):

Awesome thanks!

OpenStudy (cwrw238):

yw

OpenStudy (cwrw238):

13-12-5 is one of the 'whole number' triangles which fit the pythagoras theorem